# Quantitative U/Th deposition and cleanliness control strategies in the JUNO site air

### v1

@@ Abstract
The Jiangmen underground neutrino observatory (JUNO) is made of a 20 kt liquid scintillator (LS) detector at a depth of 700 m underground. In order to meet all physics requirements, the 238U/232Th content in the LS is required to reach a level of 10-17 g/g. However, the radioactivity of dust in the air is about 12 orders of magnitude higher than that, so there is an extremely high requirement for the cleanliness of the installation environment on site. In this study, a clean room management mode was implemented in the 120,000 m3 space of the JUNO underground experimental main hall, to control the environmental cleanliness at a level equivalent to a class 10,000-100,000 clean room. Additionally, we designed a method to directly measure the deposition rate of 238U/232Th on the surface of the detector. Based on ICP-MS detection, the sensitivity to 238U/232Th concentrations can reach the level of picograms (pg). This helps to implement the cleanliness control strategies and to assess the level of external contamination during the construction of the detector.
July 8, 2025
@@ I Introduction
Dust in the air typically contains radioactive 238U/232Th at the level of ppb-ppm. Therefore, for ultra-low background detection technology and low-background experiments, there are strict requirements for environmental cleanliness. The underground rock at the Jiangmen Underground Neutrino Observatory (JUNO) site contains 9.6 ppm of 238U and 26.6 ppm of 232Th JUNO:2021vlw . The walls of the experimental hall have only been sandblasted and not specially protected, allowing rock powder to disperse into the air and settle onto the surface of the detector during construction activities and ventilation.
@@
The main target medium of the JUNO experiment is a 20 kt liquid scintillator (LS) contained in a 35.4 m diameter acrylic sphere. To meet all JUNO physics requirements, the 238U/232Th content in the LS is required to reach a level of 10-17 g/g, which is about 12 orders of magnitude lower than the 238U/232Th content in the rock, thus imposing extremely high cleanliness requirements on the detector and its surroundings. Additionally, if too much dust accumulates on the outside of the acrylic sphere, the gamma radiation resulting from 238U/232Th decay in the dust can penetrate the water layer (see Figure 1) and contribute to the background of the LS. This necessitates stringent cleanliness requirements not only for the acrylic vessel containing the LS but also for the entire installation environment in the hall.
@@
Furthermore, in an environment which requires a certain level of cleanliness, it is necessary to quantitatively measure how much natural radioactive 238U/232Th will settle on the detector. Inductively Coupled Plasma Mass Spectrometry (ICP-MS) has extremely high sensitivity for measuring 238U/232Th, allowing for the design of experiments to quantitatively measure the actual 238U/232Th deposition rate under known levels of environmental cleanliness. This work will provide a detailed overview of the cleanliness control strategies implemented at the JUNO site and the current level of cleanliness achieved.
The structure of the article is as follows: Section II describes the detector construction in the JUNO experimental hall, Section III describes the cleanliness control indicators at the JUNO site, Section IV presents the cleanliness control strategies and the long-term cleanliness monitoring results in the hall, Section V describes the experiments and results of 238U/232Th deposition rates in the air based on ICP-MS, and Section VI presents the application to JUNO, Section VII is the summary of this work.
@@ II Overview of the JUNO detector
The conceptual design of the JUNO detector is illustrated in Figure 1. A 20 kt LS, serving as the target for neutrino detection, is contained within a 35.4-meter-diameter acrylic sphere. The entire acrylic sphere is supported by a 41.1-meter-diameter Stainless Steel (SS) truss, which is held in place by SS support rods. A total of 17,612 20-inch and 25,600 3-inch Photomultiplier Tubes (PMTs) are installed on the SS truss, designed to detect the faint light signals generated by neutrino interactions with the LS. The entire spherical structure is housed within a cylindrical cavity measuring 43.5 m in diameter and 44 m in height. This cavity is filled with 40 kt of ultra-pure water, which, in combination with the top-mounted track detector, serves to mark cosmic rays and shield against environmental background. Additionally, magnetic shielding coils are installed on the SS truss to mitigate the interference of the Earth’s magnetic field on the PMTs. A calibration room is positioned at the top of the spherical detector, approximately 9 meters above the surface of the acrylic sphere. Due to the lower density of the LS compared to water, the liquid level inside the acrylic sphere is maintained 4.5 m higher than the external water level to ensure pressure equilibrium. Consequently, a 9-meter chimney extends upward from the top of the acrylic sphere. To account for light blocking, mechanical strength, and background effects, the lower 1 m section of the chimney is made of acrylic, while the remaining portion is constructed of stainless steel. The chimney’s upper end extends into the top calibration room, allowing calibration sources to be introduced into the detector’s interior through the chimney for calibration purposes.
@@ III Cleanliness requirement
JUNO high-purity scenario requires the 238U/232Th concentration in its 20 kt LS to be at the level of 10-17 g/g, corresponding to a maximum of 0.2 g for both 238U and 232Th. After the completion of the detector construction, the JUNO acrylic sphere will first be filled with Ultra-Pure Water (UPW) from the bottom, then the LS will be injected from the top to replace the UPW inside the sphere. During the filling process, the air inside the sphere will come into direct contact with the acrylic surface and the UPW. The most conservative assumption is that all the dust in the air inside the sphere will be transferred to the LS. The dust in the air mainly comes from rock powder, with a radioactive assumption consistent with that of the rock (9.6 ppm 238U and 26.6 ppm 232Th). Therefore, it is required that the dust content in the air be less than 8 mg, dominated by 232Th concentration. According to the particle size distribution in the cleanroom, assuming that the particles are spherical and their density is similar to the rock of 2.6 g/cm3, the total particle mass in a cleanroom of class 1,000 ***A class 1,000 environment is equivalent to ISO class 6 according to ”Federal Standard 209E”. is 12 mg, which is close to the requirement for the LS. During the actual filling process, the air inside the sphere will be vented from the top, and the dust in the air will not all settle on the water and acrylic surfaces. However, we conservatively require the cleanliness of the air inside the sphere to reach class 1,000 before filling it with water.
@@
The space of the JUNO experimental main hall is approximately 120,000 m3, and it is difficult to control the cleanliness level of such a large space within the class 1,000 standard. Therefore, after the construction of the acrylic sphere, the upper and lower chimney openings were temporarily sealed with a plate to avoid significant exchange of internal and external gases. The inner surface of the acrylic sphere was protected by a film during the construction process. The main material of this film is paper, with one side coated with a thin layer of water-soluble adhesive. During the final cleaning process, the adhesive dissolves, allowing the paper film attached to the surface of the organic glass to naturally fall off. Therefore, before removing the protective film, it is possible to use 3D nozzles to generate micron-level water mist, allowing the dust in the air inside the sphere to adhere to the water droplets and settle to the bottom of the acrylic sphere, thereby further improving the air quality inside the sphere. Based on usual observations, the concentration of particulate matter in urban air significantly improves immediately after rain, typically reducing by one order of magnitude. With better optimization and design at JUNO site, we expect the effect of dust reduction through water mist in the 35.4 m diameter acrylic sphere allows an improvement of one to two orders of magnitude in the cleanliness. Therefore, we require the initial state inside the sphere – that is, in the installation environment – to reach a cleanliness level of class 10,000-100,000 in the installation environment of the hall.
The dust deposited on the outer surface of the acrylic sphere, PMT, and stainless steel (SS) truss during the construction process, can produce gamma rays by the natural radioactivity inside. The gamma rays can penetrate through the water layer and the acrylic sphere and enter the LS, constituting the background. Based on the JUNO software framework Lin:2022htc , we simulated the natural radioactivity on the outer surface of the acrylic sphere, PMT, and the SS truss JUNO:2021kxb . The probability for the deposited energy of being greater than 0.7 MeV and be positioned within the fiducial volume (R 17.2 m) of the LS for each natural radioactive nuclide is listed in Table 1. Taking the acrylic outer surface as an example, 1 Bq of 238U will lead to 0.001 Hz event rate in the LS with deposited energy larger than 0.7 MeV at a position with radius smaller than 17.2 m, while the rate is one order of magnitude larger for 232Th, which has more energetic gamma rays in the chain (208Tl). At the same time, the radon produced by 238U in the dust will enter the shielding water, causing an increase of radon concentration. The background impact of the outer detector requires knowledge of the 238U/232Th settling rate of the dust in the air. To this aim, we designed an experiment to expose containers and pads, with extremely low background levels, to a known cleanliness environment for a certain period of time. We then measured the 238U/232Th content deposited on the surface of the containers and pads using ICP-MS. The experimental details and results will be presented in detail in Section V.
Acrylic outer surface | PMTTruss | ||
Radius [m] | 17.8 | 19.2-20.0 | |
Probability | 238U | 1.110-3 | 3.410-6 |
232Th | 1.310-2 | 1.410-5 | |
40K | 1.110-3 | 1.310-7 |
@@ IV.1 Overall environment
In early 2022, the SS truss construction began in an installation environment that was still very poor at that time. The hall and ancillary rooms were not sealed and were connected to the outside traffic support tunnel. The rock walls inside the hall were only sandblasted, without any other special protection. There were four fume hoods in the hall with a circulation and filtration capacity of 200,000 m3/h to improve the cleanliness of the hall. The weathered rock powder on the surface of the rock walls would float into the air with the airflow, and the activities of the installation personnel would also bring dust. The air cleanliness was at a level of class 1,000,000.
In order to improve the cleanliness of the hall environment, we carried out a deep cleaning of the hall in April 2022. The rock walls above the ground level of 3 m were cleaned with water to reduce dust caused by personnel activities. For areas above 3 m, water flushing was challenging on one hand, and the ducts had sponges and could not be flushed with a large amount of water. However, it is believed that with long-term circulation and filtration, the dust that can be carried away by the airflow will gradually decrease. All surfaces such as the floor, handrails, and ducts were vacuumed and then wiped again. The walls of the cylindrical water pool, with a diameter of 43.5 m and height of 44 m, were rinsed with deionized water, and the installed SS structure was also rinsed with deionized water.
The cargo shower rooms at the top and bottom entrances of the hall were completed and put into operation in early May 2022. From then on, all goods and personnel entering the hall needed to pass through the air shower to remove surface dust. In the same month, fire doors were installed in the ancillary rooms around the hall, effectively achieving isolation between the hall and the outside traffic support tunnel. At the same time, we set up dressing areas at the entrances of the two cargo shower rooms to implement a clean room management mode in the hall. Personnel entering the hall needed to change into special anti-static clean suits in this area. Plastic mats were laid on the cement floor inside the hall to facilitate cleaning and reduce dust caused by personnel movement on the cement floor.
@@
All goods and equipment entering the hall must be cleaned outside the hall, wrapped in protective film during transportation, and the film must be removed at the hall entrance and passed through the air shower in the cargo shower room before entering the hall.
The paper outer packaging of all goods and equipment must be removed at the entrance of the cargo shower room before entering, and only low-dust paper boxes customized by PMT are allowed to enter.
@@
In terms of management, the overall communication and coordination are the responsibility of the head of the background task force, the heads of the various subsystems are responsible for review and implementation, and the quality inspection is carried out by the gatekeeper.
We conducted cleanliness level measurements at multiple points inside the hall and found that the differences were minimal. This is because a circulating purification fan with a capacity of 200,000 m³/h is operating in the hall, which results in a relatively uniform distribution of particulate matter in the air. Therefore, we selected a fixed point within the hall for long-term stability testing. The long time monitoring of cleanliness inside the hall is shown in Figure 2. Some key clean control measures have been marked in the figure. According to the particle size distribution in the class 10,000 environment (ISO class 7), all particles can be calculated as spherical to obtain the total particle volume per unit volume of air at the class 10,000 level. We continuously monitor the particle size distribution in the air inside the hall using a laser dust particle counter and calculate the real-time total particle volume per unit volume based on spherical particles. The values on the Y-axis in the graph represent the real-time total particle volume divided by the total particle volume at the class 10,000 level, indicating the cleanliness level relative to the class 10,000 level. With all of the efforts described above, the cleanliness inside the experimental hall has reached the equivalent of a class 10,000-100,000 level since May 2022, with an average cleanliness level of class 74,000. The 35.4 m diameter acrylic sphere is assembled on-site with 265 acrylic panels, with each layer being meticulously polished to improve its smoothness. Starting in 2023, the installation of acrylic increased the number of construction workers, and the installation schedule was changed to two shifts per day. Some of the larger peaks from March to July in the graph mainly resulted from the dispersion of acrylic powder into the air during the polishing of the acrylic panel joints. During non-polishing periods, the cleanliness level in the hall mostly remains at class 10,000. Since the acrylic material we used to build the acrylic sphere was very clean – less than 10-12 g/g measured U/Th concentrations, the acrylic powder generated during the polishing process has minimal impact on the background compared to rock dust.
@@ IV.2 Detector cleaning and protection
The SS truss installation of the JUNO detector was finished in June 2022: the truss then underwent three extensive cleanings in 2022, including water flushing and wiping processes. Moreover, before installing each of the PMT layers, we conducted another thorough cleaning of the corresponding area in the SS truss.
The installation of the acrylic sphere began in July 2022 and was completed by the summer of 2023 at the equatorial layer, with the lower hemisphere installation completed in October 2024. The outer surface of the acrylic sphere was covered with a protective film throughout the installation process. Unlike the film attached to the inner surface, the film on the outer surface is made of polyethylene with a minimal amount of regular adhesive on one side. This film requires manual removal, and laboratory tests have confirmed that the residual background impact of the adhesive is negligible. From September to October 2024, we completely removed the protective film, cleaned the acrylic outer surface using a clean cloth, and painted the seam with epoxy. The inner surface of the acrylic sphere was also protected by another type of film during the construction process, which was ultimately removed by final 3D high-pressure spraying inside the sphere.
The mass installation of PMTs started in early 2023 and reached the equatorial layer by April 2024, with the completion of the lower hemisphere installation in December 2024. In February 2024, we cleaned the surface of the installed PMT area with an electrostatic duster.

### v2

@@ Abstract
The Jiangmen Underground Neutrino Observatory (JUNO) employs a 20 kt liquid scintillator (LS) detector located 700 m underground. To meet its physics objectives, the LS must achieve an ultra-low 238U/232Th content of 10-17 g/g. Given that airborne dust exhibits radioactivity about 12 orders of magnitude higher, exceptional cleanliness is essential during on-site installation. The total permissible dust mass in the 20 kt LS is only about 8 mg. To attain this, the acrylic vessel interior must comply with class 1,000 cleanliness. Pre-filling water spray cleaning improves cleanliness by roughly two orders of magnitude, requiring the overall environment to be maintained between class 10,000 and 100,000. At JUNO, a cleanroom management system has been implemented across the 120,000 m3 underground experimental hall. Since May 2022, continuous laser particle monitoring has consistently achieved an average cleanliness class of 74,000. Furthermore, we developed a method to directly measure 238U/232Th deposition rates on detector surfaces. Using ICP-MS, sensitivity reaches sub-ppt levels (10-12 g/g), enabling effective cleanliness control and assessment of external contamination during detector construction.
@@ I Introduction
Dust in the air typically contains radioactive 238U/232Th at the level of ppb-ppm. Therefore, for ultra-low background detection technology and low-background experiments, there are strict requirements for environmental cleanliness. The underground rock at the Jiangmen Underground Neutrino Observatory (JUNO) site contains 9.6 ppm of 238U and 26.6 ppm of 232Th Abusleme and others (2022). The walls of the experimental hall have only been sandblasted and not specially protected, allowing rock powder to disperse into the air and settle onto the surface of the detector during construction activities and ventilation.
@@
The main target medium of the JUNO experiment is a 20 kt liquid scintillator (LS) contained in a 35.4-meter-diameter acrylic sphere. To meet all JUNO physics requirements, the 238U/232Th content in the LS is required to reach a level of 10-17 g/g Abusleme and others (2022), which is about 12 orders of magnitude lower than the 238U/232Th content in the rock. The dust present on the inner surface of the acrylic vessel and within the internal environment were in direct contact with the LS during the filling, thus we imposed extremely high cleanliness requirements on the acrylic sphere and its surroundings. Additionally, if too much dust accumulates on the detector components outside the acrylic sphere, such as Photomultiplier Tubes (PMTs), the Stainless Steel (SS) structure, and Ultrapure Water (UPW), the gamma radiation resulting from 238U/232Th decay in the dust can penetrate the water layer (see Fig. 1) and contribute to the background in the LS. This necessitates stringent cleanliness requirements not only for the acrylic vessel containing the LS but also for the entire installation environment in the hall.
@@
Furthermore, in an environment requiring a certain level of cleanliness, it is necessary to quantitatively measure how much natural radioactive 238U/232Th will settle on the detector. Inductively Coupled Plasma Mass Spectrometry (ICP-MS) has extremely high sensitivity for measuring 238U/232Th, allowing for the design of experiments to quantitatively measure the actual 238U/232Th deposition rate under known levels of environmental cleanliness.
This work will provide a detailed overview of the cleanliness control strategies implemented at the JUNO site and the current level of cleanliness achieved. The structure of the article is as follows: Section II presents the conceptual design of the JUNO detector. Section III explains how the radiopurity requirements for the JUNO LS were translated into cleanliness standards for the air in direct contact with the LS inside the acrylic vessel, and how these standards were extended to the entire experimental hall environment. Building on these criteria and considering actual installation conditions, Section IV describes the cleanliness control strategy developed for the installation process and evaluates its effectiveness through long-term monitoring of particulate levels in the hall. Section V presents a precise experiment conducted during the installation phase to measure the U/Th deposition on surfaces of various shapes and orientations over a defined period. This experiment enabled the quantification of expected U/Th accumulation on detector surfaces throughout the installation. Finally, based on measured U/Th deposition rates and simulations of natural radioactivity at different detector material locations, Section VI estimates the background contribution in the LS from U/Th deposited on detector surfaces during installation, as well as its impact on radon contamination levels in the external UPW region. Section VII is the summary of this work.
@@ II Overview of the JUNO detector
The conceptual design of the JUNO detector is illustrated in Fig. 1. A 20 kt LS, serving as the target for neutrino detection, is contained within a 35.4-meter-diameter acrylic sphere. The entire acrylic sphere is supported by a 41.1-meter-diameter SS truss, which is held in place by SS support rods. A total of 17,612 20-inch and 25,600 3-inch PMTs are installed on the SS truss, designed to detect the faint light signals generated by neutrino interactions with the LS. The entire spherical structure is housed within a cylindrical cavity measuring 43.5 m in diameter and 44 m in height. This cavity is filled with 40 kt of UPW, which, in combination with the top-mounted track detector, serves to tag cosmic rays and shield against environmental background. Additionally, magnetic shielding coils are installed on the SS truss to mitigate the interference of the Earth’s magnetic field on the PMTs. A calibration room is positioned at the top of the spherical detector, approximately 9 meters above the surface of the acrylic sphere. Due to the lower density of the LS compared to water, the liquid level inside the acrylic sphere is maintained 4.5 m higher than the external water level to ensure pressure equilibrium. Consequently, a 9-meter chimney extends upward from the top of the acrylic sphere. To account for light blocking, mechanical strength, and background effects, the lower 1 m section of the chimney is made of acrylic, while the remaining portion is constructed of SS. The chimney’s upper end extends into the top calibration room, allowing calibration sources to be introduced into the detector’s interior through the chimney for calibration purposes.
@@ III Cleanliness requirement
JUNO high-purity scenario requires the 238U/232Th concentration in its 20 kt LS to be at the level of 10-17 g/g Abusleme and others (2022), corresponding to a maximum of 0.2 g for both 238U and 232Th in the full detector volume.
@@
The inner surface of the acrylic sphere was protected by a film (purchased from Daio Paper Corporation 1) during the construction process. The main material of this film is paper, with one side coated with a thin layer of water-soluble adhesive. During the final cleaning process, the adhesive dissolves, allowing the paper film attached to the surface of the organic glass to naturally fall off. Therefore, before removing the protective film, it is possible to use 3D nozzles to generate micron-level water mist, allowing the dust in the air inside the sphere to adhere to the water droplets and settle to the bottom of the acrylic sphere, thereby further improving the air quality inside the sphere. Based on usual observations, the concentration of particulate matter in urban air significantly improves immediately after rain, typically reducing by one order of magnitude. With better optimization and design at JUNO site, we expect the effect of dust reduction through water mist in the 35.4-meter-diameter acrylic sphere allows an improvement of one to two orders of magnitude in the cleanliness. If the initial cleanliness level inside the sphere is maintained within Class 10,000–100,000, it can be further improved to better than Class 1,000 after the water mist treatment.
After the completion of the detector construction, the upper and lower chimney openings were temporarily sealed with a plate to avoid significant exchange of internal and external gases. The JUNO acrylic sphere will first be filled with UPW from the bottom, then the LS will be injected from the top to replace the UPW inside the sphere. During the water filling process, the air inside the sphere will come into direct contact with the acrylic surface and the UPW. The most conservative assumption is that all the dust in the air inside the sphere will be transferred to the LS. The dust in the air mainly comes from rock powder, with a radioactive assumption consistent with that of the rock (9.6 ppm 238U and 26.6 ppm 232Th). Therefore, it is required that the dust content in the air be less than 8 mg, dominated by 232Th concentration. According to the particle size distribution in the cleanroom, assuming that the particles are spherical and their density is similar to the rock of 2.6 g/cm3, the total particle mass in a cleanroom of class 1,000 (equivalent to ISO class 6 according to ”Federal Standard 209E”) and size 23,216 m3 is 12 mg, which is close to the requirement for the LS. During the actual filling process, the air inside the sphere will be vented from the top, and the dust in the air will not all settle on the water and acrylic surfaces. However, we conservatively require the cleanliness of the air inside the sphere to reach class 1,000 before filling it with water.
@@ IV.1 Overall environment
In early 2022, the SS truss construction began in an installation environment that was still very poor at that time. The hall and ancillary rooms were not sealed and were connected to the outside traffic support tunnel. The rock walls inside the hall were only sandblasted, without any other special protection. The weathered rock powder on the surface of the rock walls would float into the air with the airflow, and the activities of the installation personnel would also bring dust. Air cleanliness was measured with a particle counter, and the particle concentration is estimated to be comparable to that of Class 1,000,000. There were four self-circulating filtration fans in the hall with a circulation and filtration capacity of 200,000 m3/h to improve the cleanliness of the hall. All self-circulating filtration fans are equipped with a three-stage filtration system, consisting of primary, medium, and high-efficiency filters. However, the hall is not isolated from the external environment, and there are no clean protective measures for the movement of personnel and goods. As a result, even with four self-circulating filtration fans, the capacity for environmental improvement remains limited.
In order to improve the cleanliness of the hall environment, we carried out a deep cleaning of the hall in April 2022. The rock walls above the ground level of 3 m were cleaned with water to reduce dust caused by personnel activities. For areas above 3 m, water flushing was challenging on one hand, and the ducts had sponges, which could not be flushed with a large amount of water. However, residual dust in the experimental hall is expected to gradually diminish over time through continuous circulation and filtration. All surfaces above and below 3 m, such as the floor, handrails, and ducts were vacuumed and then wiped again. The walls of the cylindrical water pool, with a diameter of 43.5 m and height of 44 m, were rinsed with deionized water, and the installed SS structure was also rinsed with deionized water.
The cargo air shower rooms at the top and bottom entrances of the hall were completed and put into operation in early May 2022. From then on, all goods and personnel entering the hall needed to pass through the air shower to remove surface dust. In the same month, fire doors were installed in the ancillary rooms around the hall, effectively achieving isolation between the hall and the outside traffic support tunnel. At the same time, we set up dressing areas at the entrances of the two cargo air shower rooms to implement a clean room management mode in the hall. Personnel entering the hall needed to change into special anti-static clean suits in this area. Plastic mats were laid on the cement floor inside the hall to facilitate cleaning and reduce dust caused by personnel movement on the cement floor.
@@
All goods and equipment entering the hall must be cleaned outside the hall, wrapped in protective film during transportation, and the film must be removed at the hall entrance and passed through the air shower in the cargo air shower room before entering the hall.
The paper outer packaging of all goods and equipment must be removed at the entrance of the cargo air shower room before entering, and only low-dust paper boxes customized for PMT package are allowed to enter.
@@
In terms of management, the overall communication and coordination are the responsibility of the head of the background task force, the heads of the various subsystems are responsible for review and implementation, and the quality inspection is carried out by a gatekeeper.
We conducted cleanliness level measurements with a particle size analyzer (Lighthouse Handheld 3016) at multiple points inside the hall and found that the differences were within 10%. This is primarily due to the continuous operation of four circulating air purification fans with a capacity of 200,000 m3/h, throughout the entire installation period. Their sustained operation promotes effective mixing and circulation of air, leading to a relatively uniform distribution of particulate matter in the hall. Therefore, we selected a fixed point within the hall for long-term stability testing. The long time monitoring of cleanliness inside the hall is shown in Fig. 2. Some key clean control measures have been marked in the figure. According to the particle size distribution in the class 10,000 environment (ISO class 7), all particles can be considered as spherical to obtain the total particle volume per unit volume of air at the class 10,000 level. We continuously monitor the particle size distribution in the air inside the hall using the particle size analyzer and calculate the real-time total particle volume per unit volume based on spherical particles. The values on the Y-axis in the graph represent the real-time total particle volume divided by the total particle volume at the class 10,000 level, indicating the cleanliness level relative to the class 10,000 level. With all of the efforts described above, the cleanliness inside the experimental hall has reached the equivalent of a class 10,000-100,000 level since May 2022, with an average cleanliness level of class 74,000. The above cleanliness control measures have proven highly effective and can be applied to contamination prevention in other experiments of the similar condition.
The 35.4-meter-diameter acrylic sphere is assembled on-site with 265 acrylic panels, with each layer being meticulously polished to improve its smoothness. Starting in 2023, the installation of acrylic increased the number of construction workers, and the installation schedule was changed to two shifts per day. Some of the larger peaks from March to July in the graph mainly resulted from the dispersion of acrylic powder into the air during the polishing of the acrylic panel joints. During non-polishing periods, the cleanliness level in the hall mostly remains at class 10,000. Since the acrylic material we used to build the acrylic sphere was very clean – less than 10-12 g/g measured U/Th concentrations, the acrylic powder generated during the polishing process has minimal impact on the background compared to rock dust.
@@ IV.2 Detector cleaning and protection
The SS truss installation of the JUNO detector was finished in June 2022: the truss then underwent three extensive cleanings in 2022, including water flushing and wiping processes. Moreover, before installing each of the PMTs layers, we conducted another thorough cleaning of the corresponding area in the SS truss.
The installation of the acrylic sphere began in July 2022 and was completed by the summer of 2023 at the equatorial layer, with the lower hemisphere installation completed in October 2024. The outer surface of the acrylic sphere was covered with a protective film throughout the installation process. Unlike the film attached to the inner surface, the film on the outer surface is made of polyethylene with a minimal amount of regular adhesive on one side. This film requires manual removal, and laboratory tests have confirmed that the residual background impact of the adhesive is negligible Li and others (2023). From September to October 2024, we completely removed the protective film, cleaned the acrylic outer surface using a clean cloth, and painted the seam with epoxy (radiopurity screened by ICP-MS: 5 ppt 238U and 7 ppt 232Th). The inner surface of the acrylic sphere was also protected by another type of film during the construction process, which was ultimately removed by final 3D high-pressure spraying inside the sphere.
The mass installation of PMTs started in early 2023 and reached the equatorial layer by April 2024, with the completion of the lower hemisphere installation in December 2024. In February 2024, we cleaned the surface of the installed PMTs area with an electrostatic duster.